OpenStudy (ellie202000):

Use the Point-Slope Form to write the equation of lines P, Q, and R. Then, write the equation in slope-intercept form.

OpenStudy (jdoe0001):

P,Q,R? they kinda went MIA there

OpenStudy (ellie202000):

They give no graph of coordinates. I don't know where to start.

OpenStudy (jdoe0001):

well, they're referring to an equation for those segments...so, check your material, should be there, or email your teacher asking if it's somewhere where it might be

OpenStudy (ellie202000):

I know that point-slope form is \[y-y1=m(x-x1)\] So for example. One set of coordinates is (0,0) and the slope is 3/2 the equation would be \[y-0=\frac{ 3 }{ 2 }(x-0)\] which equals \[y=\frac{ 3 }{ 2 }x\]

OpenStudy (ellie202000):

but I don't what to do with this problem.

OpenStudy (ellie202000):

@dan815 @DanJS @undeadknight26

OpenStudy (jdoe0001):

@Ellie202000 either ask your teacher by email or phone, where to get P, Q and R equations or you can always try the 1-900 of the psychic hotline may also want to ask them to give you a palm reading as well =)

Vocaloid (vocaloid):

no other information is given? then it would be impossible because we don't know what the lines P, Q, and R look like

OpenStudy (ellie202000):

Thank you guys for looking at taking the time for my question. Appreciate it. :)

OpenStudy (danjs):

WIth no information, you probably just have to use a general form with all the variables \[(P _{x1}, P_{y1})\] \[(P _{x2}, P_{y2})\] as two points for line P for example